OpenStudy (mathiscancer):

So I have a right triangle, all i'm given is the Hypotenuse and an angle. The Hypotenuse is 10, I need to find X, the Adjacent. I figured out all the angles and they are the following: 90, 36(given), and 54. I really can't figure this out and i need some guidance as i really don't understand cos, sin, and tan... at all.

OpenStudy (kinged):

Use law of sines.

OpenStudy (kinged):

(The sin of an angle)/(the length of the opposite sides) is the same for all the angles.

OpenStudy (mathiscancer):

How would i find the sin of the angle, and which angle?

OpenStudy (kinged):

Use the calculator to get the sin of an angle.

OpenStudy (mathiscancer):

that's not a right triangle, does that matter?

OpenStudy (kinged):

It is right.

OpenStudy (mathiscancer):

that picture you sent me isn't a right triangle

OpenStudy (kinged):

Oh. It works for all triangles. Thought you were talking about your triangle.

OpenStudy (mathiscancer):

Oh, let me see if i can figure it out real quick...

OpenStudy (mathiscancer):

Wait, so how exactly do I plug in the sides if I don't have them all?

OpenStudy (kinged):

You only need one.

OpenStudy (kinged):

The ratio is the same for all three angles. You're given three angles and a side.

OpenStudy (mathiscancer):

sooo... Sin(54)|10 = Sin(36)|10 = Sin(90)|10

OpenStudy (mathiscancer):

Is that right?

OpenStudy (kinged):

Oh no.

OpenStudy (jh99):

SOH CAH TOA

OpenStudy (mathiscancer):

So then how would I do this?

OpenStudy (kinged):

\[\frac{ \sin90 }{ 10 }=\frac{ \sin36 }{ opposite(side) }\]Solve for (opposite side).

OpenStudy (kinged):

That's just for the first unknown side.

OpenStudy (kinged):

We'll do the same for the last unknown side.